if h(r)= 2/3r-6, what is the value of h(-9)?
Answer by nerdybill(7384)

if h(r)= 2/3r-6, what is the value of h(-9)?
.
substitute -9 for r and solve:
h(r)= 2/3r-6
h(-9)= 2/3(-9)-6
h(-9)= 2(-3)-6
h(-9)= (-6)-6
h(-9)= -12